# Surface decomposition method for sensitivity analysis of first-passage dynamic reliability of linear systems

### v1

@@ 5 Conclusions
A key advantage of the surface decomposition method lies in its ability to reuse function evaluation results for sensitivity analyses with respect to different design parameters, thereby substantially reducing the computational cost when a large number of parameters are considered. The effectiveness of the proposed method has been validated by two first-passage dynamic reliability problems of linear systems. It has been found that, for the numerical examples studied, only less than 2000 function evaluations are required to obtain the sensitivity estimates of the first-passage failure probability with respect to all design parameters with a target coefficient of variation of 0.1. Furthermore, similar to the efficient importance sampling and the domain decomposition method for first-passage dynamic reliability analysis, the surface decomposition method for reliability sensitivity analysis also demonstrates a higher efficiency as the failure probability becomes smaller. The proposed method exhibits great potential for reliability-based parametric and topology optimization.

### v2

@@ 5 Conclusions
The effectiveness of the surface decomposition method has been validated by three first-passage dynamic reliability problems of linear systems, one of which involves a large-scale building frame structure. It has been found that, for the numerical examples studied, only less than 2000 function evaluations are required to obtain the sensitivity estimates of the first-passage failure probability with respect to all the design parameters with a target coefficient of variation of 0.1. The surface decomposition method exhibits a higher efficiency than the directional importance sampling with a speed-up ratio ranging from 1.29 to 2.56. Furthermore, similar to the efficient importance sampling and the domain decomposition method for first-passage dynamic reliability analysis, the surface decomposition method for reliability sensitivity analysis also demonstrates a higher efficiency as the failure probability becomes smaller. Numerical studies also indicate that the efficiency of the proposed method is not affected by the input dimensionality, while it decreases to a certain degree as the number of component failure events increases.
The core contribution of this work is that it represents the first attempt to explicitly address the complex surface integral arising in reliability sensitivity analysis, which has long been regarded as highly challenging and thus largely bypassed in the literature. By introducing a novel yet simple concept of surface decomposition, the proposed method allows efficient evaluation of this integral for linear systems under Gaussian excitations and exhibits favorable performance compared with existing state-of-the-art approaches, while relying on an algorithm that is straightforward to implement and readily reproducible. A key advantage of the proposed method is its ability to reuse the function evaluation results for sensitivity analyses with respect to different design parameters, thereby substantially reducing the computational cost when a large number of parameters are considered. This feature highlights the great potential of the method for reliability-based parametric and topology optimization.
The current surface decomposition method has two main limitations: system linearity and excitation Gaussianity. The root cause of these limitations is that the efficient implementation of the method relies on the availability of closed-form expressions for dynamic responses and response sensitivities at different time instants. Without the assumptions of linear system behavior and Gaussian excitation, such closed-form expressions are either unavailable or significantly more difficult to obtain. The first limitation can be mitigated by incorporating the equivalent linearization technique, although this approach is generally restricted to mildly nonlinear systems and inevitably entails some loss of accuracy in reliability sensitivity estimation. The second limitation can be addressed for certain types of non-Gaussian excitations with specific forms, such as the quadratic Gaussian excitations, for which the explicit expressions of dynamic responses and response sensitivities can still be derived in quadratic form. A method to address this limitation is currently under investigation by the authors.